What is the mass of iron if 842.05 joules are added and the temperature moves from 25 oC to 28 oC?

sFe = 0.449 J/(gK)


1
Expert's answer
2021-03-08T06:14:23-0500

According to the heat equation:

Q = mc(T2-T1)

where Q - absorbed or released heat (842.05 J), c - specific heat (0.449 J g-1 K-1), T2 - final temperature (28 oC = 301.15 K), T1 - initial temperature (25 oC = 298.15 K).

From here:

mFe = Q / cFe(T2-T1) = 842.05 J / [0.449 J g-1 K-1 × (301.15 K - 298.15 K)] = 625.13 g


Answer: 625.13 g